# Tailoring transport in quantum spin chains via disorder and collisions

### v3

@@ III.3 Entanglement Entropy
To complement the studies based on local observables, in this section we analyze the behavior of entanglement as a measure for thermalization dynamics in our spin chain. We focus on the case of two excitations injected into the system next to each other, that is the most interesting situation found in previous sections, where disorder, noise and coupling between spins significantly affected the delocalization behavior. We limit to two excitations for computational ease, having now clear what does change from two to more. Finally, we again zoom the parameters in the interesting range emerged in Sec. III.2.1, that allows the emergence of plateaus and non-continuous temporal profiles of transport: this is low collision rates, e.g. with $r_{c}=0.1$, large disorder range e.g. with $h=10$ and important anisotropy e.g. with $\Delta=2.5$. Finally, we now want to check the dependence on the noise space-homogeneity by letting the shape parameter $\nu$ vary from absence of noise ($\nu=0$), heterogeneous collisions over time ($\nu\leq 1$) and time-homogeneous collisions ($\nu=100$).
@@
We then evaluate the Von Neumann entropy of the corresponding reduced density and present in Fig. 5 the time evolution of the entanglement entropy $S_{vN}(t)$ in log scale.
@@
Figure 5: Entanglement Entropy and dependence on noise space heterogeneity. $S_{vN}$ vs time $tJ$ in a log scale for different levels of collision-time homogeneity $\nu$ and fixed low collision rate $r_{c}=0.1$, large disorder range $h=10$ and anisotropy $\Delta=2.5$ in the interesting plateaus regime, for a disordered spin chain of $N=20$ sites and two injected excitations starting next to each other. Inset: same in the main figure, but in linear scale. The plot shows the mean $S_{vN}$ over 25 of the total $M=250$ trajectories.
@@
In general, we expect that the low density of excitations limits the number of degrees of freedom available for generating entanglement: fewer interacting particles results in less correlations. Furthermore, we have moderate transport of information due to the presence of disorder, anisotropy and noise, which limits the spread of quantum correlations throughout the system. In the absence of noise $\nu=0$, we observe in Fig. 5 a flat line over time, apart from some fluctuations. In fact, without noise driving perturbations that can diffuse correlations over the spin chain, the system persists in a coherent and localized state dominated only by the effects of disorder and anisotropy. For time-heterogeneous noise (low values of $\nu$) the behavior of the entanglement entropy $S_{vN}(t)$ is logarithmic, with a slow growth consistent with the breaking of a MBL-like regime with diffusive transport. Even with increasing collision time-homogeneous stroboscopic noise ($\nu=100$), we observe the ability to also disrupt localization with the formation of plateaus. While delocalization is slower, we can tailor its rate by modifying the timescales of the plateaus as discussed in the previous sections. These plateaus are clearly visible in the inset picture in Fig. 5, in linear timescale in a reduced time range.
@@ IV Discussion
In this study, we analyze how coherent and dissipative couplings can manipulate the transport properties and thermalization of a noisy, disordered XXZ spin chain using stochastic collisional noise to simulate the interactions with its environment. We identify relevant regimes in the rich landscape determined by the interplay between disorder, anisotropy, and collisional noise to shape the localization/delocalization transitions and the specific transport rates.
@@
From the simplest one excitation case, see Sec. III.1, we found that disorder tends to suppress transport through interference, which competes with collisional noise while the latter introduces dephasing, thus facilitating delocalization, instead. One first relevant message emerging from our study is that among these competing mechanisms, low collision rates ($r_{c}\ll 1$) favor the occurrence of localized regions in the form of Inverse Participation Ratio (IPR) plateaus in time windows where collisions are absent. Interestingly, the length and rate of these plateaus depend directly on noise parameters with a universal scaling, allowing to systematically tailor sequences of transport and localization that can be applied for example to stroboscopic protocols in quantum technologies. At higher collision rates ($r_{c}\geq 1$) instead, frequent collisional events completely disrupt localization, driving the system toward a diffusive regime characterized by smooth evolution of IPR over time, even for moderate levels of disorder range.
From the case of several excitations, see Sec. III.2 and Appendix B, we learn that additional features emerge due to interactions among excitations. In particular, more localized regions in the form of plateaus appear even for the tiniest levels of disorder range.
However, we still find that increasing collision rates ($r_{c}>1$) disrupt localized states, promoting transport. We demonstrate the role of noise in overcoming pinning effects, down to complete delocalization in a short time $\tau$, similar to what we found in the absence of disorder [16], where the complete delocalization times did not exceed $tJ=5$.
@@
In addition, we also find regimes of tunable transport and the presence of plateaus, in fact even at higher densities. In regimes with low collision rates, high shape parameter, and high disorder, entanglement entropy exhibits slow logarithmic growth, suggesting many-body localization behavior. This is consistent with the presence of localized states that prevent rapid spread of information. Instead, for collisions heterogeneous in time, entanglement entropy more rapidly grows. Thus, homogeneous noise leads to smoother entropy dynamics.
@@
Our analysis is of relevance for quantum technologies applications. For instance, in quantum circuits, where the efficiency of information flow directly impacts computational performance, optimizing transport is fundamental [47]. In addition, energy devices like quantum batteries and quantum heat engines could benefit from noise-engineering in order to enhance their performance in terms of efficiency and stability [23, 24]. Also, the precise manipulation of spin transport in systems such as NV centers in diamond has already enabled ultra-sensitive magnetic field measurements for quantum sensors and processors [48, 49]. Furthermore, understanding how localization-delocalization transitions arise in the presence of noise could help designing novel materials with tailored transport properties [50].
@@
Noise-disorder engineering could be used to optimize energy transport in artificial photosynthetic complexes which could use controlled noise to enhance excitonic energy transfer, mimicking the noise-assisted transport observed in natural photosynthesis. In addition, in biological systems, where energy transfer is influenced by disorder and environmental fluctuations, tailoring noise properties at microscopic level could help to balance localization and delocalization effects, aiming at maximizing efficiency. This offers means to control energy flow in quantum networks, by tuning the connectivity and interaction strength between nodes. Along these lines, the development of programmable quantum devices could be facilitated where energy or quantum information is routed dynamically.
@@ Acknowledgements.
The authors would like to thank A. Daley, G. M. Cicchini, M. C. Morrone, L. Guidi, C. Ceccanti, M.Landi and C.V. Stanzione for inspiring and useful discussions. V.S. and M.L.C. acknowledge support from the National Centre on HPC, Big Data and Quantum Computing - SPOKE 10 (Quantum Computing) and received funding from the European Union Next-GenerationEU - National Recovery and Resilience Plan (NRRP) – MISSION 4 COMPONENT 2, INVESTMENT N. 1.4 – CUP N. I53C22000690001. J.Y.M. and M.L.C. were supported by the European Social Fund REACT EU through the Italian national program PON 2014-2020, DM MUR 1062/2021. M.L.C. also acknowledges support from the project PRA2022202398 “IMAGINATION”.
Appendix A Appendix A. Imbalance
@@
In addition to IPR and IER, in the presence of disorder, the imbalance $\mathcal{I}(t)$ provides a descriptive figure of merit of the system’s relaxation dynamics. Indeed, we can write this figure of merit as:
$I=n_{e}-\tilde{n}_{o},$ (5)
where $\tilde{n}_{o}=n_{o}/N-1$, N is the number of sites and $n_{e}(n_{o})$ is the density of even (odd) particles.
@@
This definition incorporates the requirement that at time $t=0$ the excitations are only present on site(s) $n_{e}$ and imbalance $I=1$. After letting the system evolve, the excitation will spread through the system to a certain degree depending on the system parameters. In a delocalized regime we have $n_{e}=1/M$ and $n_{o}\sim 1/M$, leading to $I=0$.
In Fig. 6, we analyze the time evolution of the imbalance $\mathcal{I}(t)$ for a disordered spin chain with $N=41$ sites, one excitation starting in the middle of the chain and shape parameter $\nu=100$. In particular, Fig. 6(a) shows how the Imbalance behavior varies over time for different collision rates while Fig. 6(b) shows how it varies for different disorder range constants at fixed $r_{c}=1$. The imbalance $\mathcal{I}(t)$ quantifies the difference in population between the initially occupied and unoccupied regions of the chain, providing insight into the dynamics of the localization and delocalization of the system. For low collision rates ($r_{c}\ll 1$), the imbalance decays very slowly over time, reflecting strong localization induced by the combined effects of disorder and anisotropy.
As the collision rate increases ($r_{c}\sim 1$), noise events become more frequent, leading to a faster decay of $\mathcal{I}(t)$. This behavior is indicative of an improved delocalization, where excitation spreads more uniformly across the chain, erasing the initial population imbalance. The results highlight the critical role of collision rates in modulating transport properties: low rates stabilize localized states, while higher rates facilitate noise-assisted transport.
@@
Figure 6: Imbalance for the One excitation case. (a) Imbalance for a spin chain of $N=41$ sites and $N_{exc}=1$ excitations starting in the middle of the chain, shape parameter $\nu=100$, disorder range constant $h=10$ and interaction strength $\Delta=0$ for different collision rates. Final time is $tJ=30$. (b) Imbalance versus time for a spin chain of $N=41$ sites, one excitation, different disorder range constants at fixed collision rate $r_{c}=0.1$, shape parameter $\nu=100$ and anisotropy $\Delta=0$. The simulations were performed with $M=500$ trajectories and $dt=0.02$.
Appendix B Appendix B. Two excitations next to each other
@@
Here, we show additional results for the case of two excitations injected into the disordered spin chain. In particular, we present here what happens when anisotropy plays a more fundamental role in the dynamics of the system, by analyzing the Inverse Ergodicity Ratio (IER) under varying disorder regimes and excitation configurations. For excitations starting next to each other (Fig. 7(b)), anisotropy amplifies the formation of plateaus in the IER, even for low disorder levels. In contrast, when the excitations are initially separated (Fig. 7(a)), the effects of anisotropy are subdominant, with the dynamics primarily governed by disorder and noise. These results reinforce the particular interplay between noise, disorder, and anisotropy in modulating transport properties, offering further insights into the mechanisms driving localization and delocalization transitions. The plots are for a spin chain of $N=20$ sites, at fixed $r_{c}=0.1$, $\nu=100$, $\Delta=2.5$, varying the disorder range constants between 0.1 (low disorder) and 10 (high disorder).
@@
Figure 7: Two excitations case. Inverse Ergodicity ratio (IER) for a spin chain of $N=20$ sites and final time of simulations $tJ=30$. (a) IER vs time $tJ$ for different disorder range constants at fixed $r_{c}=0.1$, $\nu=100$, $\Delta=2.5$ for two excitations starting initially separated by two spins and (b) two excitations starting next to each other. The simulations were performed with $M=250$ trajectories and $dt=0.02$.

### v5

@@ III.3 Entanglement Entropy
To complement the studies based on local observables, in this section we analyze the behavior of entanglement as a global measure of thermalization dynamics in our spin chain. We focus on the case of two excitations injected into the system next to each other, where disorder, noise and coupling between spins significantly affect the delocalization behavior. We limit to two excitations for computational ease, having confirmed that the multiple excitations case has the same trend as the two excitations case in Sec. III.2. Finally, we again zoom the parameters in the interesting range emerged in Sec. III.2.1 with the emergence of plateaus and non-continuous temporal profiles of transport: this corresponds to low collision rates, e.g. with $r_{c}=0.1$, large disorder range e.g. with $h=10$, and important anisotropy e.g. with $\Delta=2.5$. Finally, we now want to check the dependence on the noise space-homogeneity by letting the shape parameter $\nu$ vary from absence of noise ($\nu=0$), to heterogeneous collisions over time ($\nu\leq 1$) and time-homogeneous collisions ($\nu=100$).
@@
We then evaluate the Von Neumann entropy of the corresponding reduced density and present in Fig. 7 the time evolution of the entanglement entropy $S_{vN}(t)$ with the bipartition 10-19 in a semi-log scale.
@@
Figure 7: Entanglement Entropy and dependence on noise space heterogeneity. $S_{vN}$ vs time $tJ$ in a semi-log scale for different levels of collision-time homogeneity $\nu$ and fixed low collision rate $r_{c}=0.1$, large disorder range $h=10$ and anisotropy $\Delta=2.5$ in the interesting plateaus regime, for a disordered spin chain of $N=20$ sites and two injected excitations starting next to each other. The simulations were performed with $M=250$ trajectories and timestep $dt=0.02$.
@@
In general, we expect that the low density of excitations limits the number of degrees of freedom available for generating entanglement: fewer interacting particles result in reduced correlations. Furthermore, we have moderate transport of information due to the presence of disorder, anisotropy and noise limiting the spread of quantum correlations throughout the system. In the absence of noise $\nu=0$, we observe in Fig. 7 a flat line over time, apart from some fluctuations. In fact, without noise driving the diffusion of correlations over the spin chain, the system persists in a coherent and localized state dominated only by the effects of disorder and anisotropy. For values of $\nu\neq 0$, we observe a slow logarithmic growth of $S_{vN}$, consistent with the breaking of localization due to collisions [Abanin2019, Wybo, Lukin, Fischer]. In the homogeneous noise regime ($\nu=100$), the system exhibits a slow growth between successive plateaus. Instead, when $\nu$ is smaller, we see an accelerated breaking, resulting in a faster growth of $S_{vN}$. This is due to the introduction of more collisions, being the effective collision rate $r_{c}$ higher in this case. Even with increasing collision time-homogeneous stroboscopic noise ($\nu=100$), we observe the ability to also disrupt localization with the formation of plateaus. While delocalization is slower, we can tailor its rate by modifying the timescales of the plateaus as discussed in the previous sections.
@@ IV Discussion
In this study, we analyze how coherent and dissipative couplings can manipulate the transport properties and thermalization of a noisy, disordered XXZ spin chain, where stochastic collisional noise simulates the interactions with its environment. We identify relevant regimes in the rich landscape determined by the interplay between disorder, anisotropy, and collisional noise to shape the localization/delocalization transitions and the specific transport rates.
The impact of this work is twofold. Firstly, in analyzing our datasets we could extract forms of scaling behaviors, as e.g. for the appearance of plateaus (see Figs.  3 and  4): in so doing, we make order in the plethora of data emerging from our systematic study. These scaling behaviors represent valuable guidelines to think of the system. Further studies under more complicated yet interesting conditions, such as non-conserved magnetization can follow, having our present work as a benchmark. Secondly, we have worked our results out using the paradigmatic XXZ model as playground and an enough general, two-parameters, stochastic model to shape the collisions besides the more conventional addition of disorder: in this perspective, our findings can have many potential applications in specific systems, qualitatively and quantitatively.
@@
Turning to the specific findings, from the simplest one excitation case, see Sec. III.1, we found that disorder tends to suppress transport through interference: in so doing, it competes with collisional noise, that instead facilitates delocalization through dephasing. One first message emerging from our study is that among these competing mechanisms, low collision rates ($r_{c}\ll 1$) favor the occurrence of localized regions in the form of plateaus for the Inverse Participation Ratio (IPR), in time windows where collisions are absent. Interestingly, the area of these plateaus depend directly on noise parameters with a form of universal scaling towards a saturating value. The degree of universality for this behavior certainly requires further investigation. Overall, the occurrence of these plateaus allows to systematically tailor sequences of transport and localization that can be applied for example to stroboscopic protocols in quantum technologies. At higher collision rates ($r_{c}\geq 1$) instead, frequent collisional events completely disrupt localization, driving the system toward a diffusive regime characterized by smooth evolution of IPR over time, even for moderate levels of disorder range. The presence of these localization/delocalization transitions can be understood for single excitations, since collisions do not alter the total magnetization and the system remains localized between collisions, preventing the IPR from decaying during these intervals.
@@
From the case of several excitations, see Sec. III.2 and Appendix B, we learn that additional features emerge due to interactions among excitations. In particular, more localized regions in the form of plateaus appear even for the tiniest levels of disorder range. However, we still find that increasing collision rates ($r_{c}>1$) disrupt localized states, promoting transport. We demonstrate the role of noise in overcoming pinning effects, down to complete delocalization in a short time $\tau$. This is similar to what we found in the absence of disorder [AlessandroetAl], where the complete delocalization times did not exceed $tJ=5$. In addition, we also find regimes of tunable transport and the presence of plateaus, in fact even at higher densities. In regimes of high disorder and anisotropy, collisions induce a gradual logarithmic entanglement growth. This can be understood as a signature of the progressive breaking of the localized regime, rather than its preservation, consistent with the presence of localized states that prevent rapid spread of information. For collisions heterogeneous in time instead, the entanglement entropy more rapidly grows. Thus, homogeneous noise leads to smoother entropy dynamics.
@@
Our analysis is of relevance for quantum technologies applications. For instance, in quantum circuits optimizing transport is fundamental, for the efficiency of information flow directly impacts computational performance [Quantum_circuit]. In fact, besides using thermodynamic benchmarks such as ergotropy, power output, and efficiency, energy devices like quantum batteries and quantum heat engines could benefit from noise-engineering to enhance their performance in terms of efficiency and stability [Carrega2020, Mayo2022]. Also, the precise manipulation of spin transport in systems such as NV centers in diamond has already enabled ultra-sensitive magnetic field measurements for quantum sensors and processors [Jelezko2012, Oberg2019]. In addition, understanding how localization-delocalization transitions arise in the presence of noise could help designing novel materials with tailored transport properties [Nanowires].
@@
Noise-disorder engineering could also be used to optimize energy transport in artificial photosynthetic complexes, in essence using controlled noise to enhance excitonic energy transfer, and mimicking the noise-assisted transport observed in natural photosynthesis. In addition, tailoring noise properties at a microscopic scale in biological systems, where energy transfer is influenced by disorder and environmental fluctuations, could help to balance localization and delocalization effects, aiming at maximizing efficiency. Under an alternative perspective, this offers means to control energy flow in quantum networks, by tuning the connectivity and interaction strength between nodes. Along these lines, the development of programmable quantum devices could be facilitated where energy or quantum information is routed dynamically. In fact, recent works apply reinforcement learning and neural-network optimization to tailor quantum channels for target tasks and evolutionary algorithms can optimize control pulses for population transfer in multilevel systems. These techniques can be used to design collisional sequences that maximize transport efficiency or plateau tunability [Machine_Learn_25].
Finally, one more interesting perspective is to move beyond purely Markovian collisional models, considering that time-correlated environments can qualitatively alter transport. For instance, Random Telegraph Noise (RTN) and Ornstein–Uhlenbeck Noise (OUN) preserve coherence bursts that enable high-fidelity state transfer even on complex graphs [No_markov_2025]. Embedding such correlated noise kernels into our collision model could reveal richer localization–delocalization transitions.
@@ Acknowledgements.
The authors would like to thank A. Daley, G. M. Cicchini, M. C. Morrone, L. Guidi, C. Ceccanti, M.Landi, C.V. Stanzione and S. Ausilio for inspiring and useful discussions. V.S. and M.L.C. acknowledge support from the National Centre on HPC, Big Data and Quantum Computing - SPOKE 10 (Quantum Computing) and received funding from the European Union Next-GenerationEU - National Recovery and Resilience Plan (NRRP) MISSION 4 COMPONENT 2, INVESTMENT N. 1.4 CUP - N. I53C22000690001. J.Y.M. and M.L.C. were supported by the European Social Fund REACT EU through the Italian national program PON 2014-2020, DM MUR 1062/2021. M.L.C. also acknowledges support from the project PRA2022202398 “IMAGINATION”.
Appendix A Imbalance
@@
In addition to IPR and IER, in the presence of disorder, the imbalance IMB provides a descriptive figure of merit of the system’s relaxation dynamics. Indeed, we can write this figure of merit as:
$IMB=n_{c}(t)-\tilde{n}(t),$ (5)
where $n_{c}(t)=1/2\big(1+\langle\sigma_{c}^{z}(t)\rangle\big)$ is the local density of excitations at the central site of the chain, and $\tilde{n}(t)={n(t)}/{N}$, with $n(t)=\sum_{i=1}^{N}1/2\big(1+\langle\sigma_{i}^{z}(t)\rangle\big)$ , and $N$ the total number of sites.
@@
This definition incorporates the requirement that at time $t=0$ the excitations are only present on the central site, $n_{c}(0)=1$, so that $\tilde{n}(0)=N^{-1}$ and imbalance $IMB=1-N^{-1}$. After letting the system evolve, the excitation will spread through the system to a certain degree depending on the parameters. In a fully delocalized regime we have $n_{c}=1/N$ and $\tilde{n}=1/N$, leading to $IMB=0$.
@@
Figure 8: Imbalance for the One excitation case. (a) Imbalance for a spin chain of $N=41$ sites and $N_{exc}=1$ excitations starting in the middle of the chain, shape parameter $\nu=100$, disorder range constant $h=10$ and interaction strength $\Delta=0$ for different collision rates. Final time is $tJ=30$. (b) Imbalance versus time for a spin chain of $N=41$ sites, one excitation, different disorder range constants at fixed collision rate $r_{c}=0.1$, shape parameter $\nu=100$ and anisotropy $\Delta=0$. The simulations were performed with $M=500$ trajectories and timestep $dt=0.02$.
In Fig. 8, we analyze the time evolution of the imbalance $IMB$ for a disordered spin chain with $N=41$ sites, one excitation initially localized at the middle of the chain, and shape parameter $\nu=100$. In particular, Fig. 8(a) shows how the Imbalance behavior varies over time for different collision rates, while Fig. 8(b) shows its dependence on different disorder strengths at fixed $r_{c}=1$. The imbalance $IMB$ quantifies the difference in population between the initially occupied and unoccupied regions of the chain, providing insight into the dynamics of localization and delocalization. For low collision rates ($r_{c}\ll 1$), the imbalance decays very slowly over time, reflecting strong localization induced by the combined effects of disorder and anisotropy.
As the collision rate increases, noise events become more frequent, leading to a faster decay of $IMB$. This behavior indicates improved delocalization, where excitation spreads more uniformly across the chain, erasing the initial population imbalance. However, as with the IPR, higher values of $r_{c}$ (50, 100) result in a slowdown of the dynamics of the system, manifestation of the quantum Zeno effect [AlessandroetAl].
Appendix B More data on the two excitations case
@@
Here, we show additional results for the case of two excitations injected into the disordered spin chain. In particular, we present what happens when anisotropy plays a more fundamental role in the dynamics of the system, by analyzing the Inverse Ergodicity Ratio (IER) under varying disorder regimes and excitation configurations. For excitations starting next to each other (Fig. 9(b)), anisotropy amplifies the formation of plateaus in the IER, even for low disorder levels. This can be seen as a more localized behavior in the presence of a major potential energy due to the proximity of the excitations, leading to a minor kinetic energy. In contrast, when the excitations are initially separated (Fig. 9(a)), the effects of anisotropy are subdominant, with the dynamics primarily governed by disorder and noise. These results reinforce the particular interplay between noise, disorder, and anisotropy in modulating transport properties, offering further insights into the mechanisms driving localization and delocalization transitions. The plots are for a spin chain of $N=20$ sites, at fixed $r_{c}=0.1$, $\nu=100$, $\Delta=2.5$, varying the disorder range constants between 0.1 (low disorder) and 10 (high disorder). We report here also a semi-log 2D version of the 3D plot in Sec. III.2 Fig. 4. Focusing on higher values of the disorder range constant $h=5,10$, which exhibit plateaus, the trend of $D$ is $1/r_{c}$ as happened in the case of the 3D plot. Finally, in Fig. 4(b) we show a 2D plot of the complete delocalization time $\tau$ as a function of the collision rate $r_{c}$ of Fig. 5, where the shape parameter $\nu$ and the disorder strength vary as indicated in the legend. In agreement with the 3D version, it is possible to recognize the two regions: the one without the plateaus (1) with small values of $\tau$; and the one with the formation of plateaus corresponding to the regimes (2) and (3), which has higher values of $\tau$. In addition, it is possible to see here that, for $r_{c}>5$, as the system-environment interactions become more frequent, the dynamics is slowed down, leading to a gradual increase of the complete delocalization time $\tau$. This is a manifestation of the Zeno effect, also discussed in Sec. III.1 and App. A.
@@
Figure 9: Two excitations case. Inverse Ergodicity ratio (IER) for a spin chain of $N=20$ sites and final time of simulations $tJ=30$. (a) IER vs time $tJ$ for different disorder range constants at fixed $r_{c}=0.1$, $\nu=100$, $\Delta=2.5$ for two excitations starting initially separated by two spins and (b) two excitations starting next to each other. The simulations were performed with $M=250$ trajectories and timestep $dt=0.02$.
Figure 10: 2D versions of the plots in Sec. III.2. (a) A semi-log version of the Duration $D$ of the plateaus in terms of tJ vs the product $tJ\ r_{c}$, with $h=5,10$ varying the shape parameter as in the legend. (b) Complete delocalization time $\tau$ in Fig. 5 vs the collision rate $r_{c}$. The shape parameter $\nu$ and the disorder strength $h$ vary with the color and shape legend, respectively. Also in this 2D version it is possible to recognize the two regions: without the plateaus with small values of $\tau$ (1); with the formation of plateaus corresponding to the regimes (2) and (3), which has higher values of $\tau$. The simulations were performed for a spin chain of $N=20$ sites with two excitations initially separated by two spins, at fixed anisotropy $\Delta=2.5$, with $M=250$ trajectories and timestep $dt=0.02$. Final time of the simulations is $tJ=30$